The second angle of a triangle is three times as large as the first. The third angle is 20 degrees less than the sum of the other two angles. Find the measure of the first angle.
Answer by tomlaidlaw(14)

x + 3x + 4x-20 = 180
8x - 20 = 180
8x = 200
x = 25
3x = 75
4x - 20 = 80